OpenStudy (anonymous):

having trouble with graphing 7x − 9y ≥ 12

OpenStudy (anonymous):

my answer is y≤ 7/9x + 4/3

OpenStudy (anonymous):

(7x-12)/9≥y

OpenStudy (anonymous):

The way I learned it was to use the slope to find a point using a trick that my teacher taught me to remember: Ryse/Run so for an equation such as y=7/9X I would have 0,0 and then the point 9,7. Since 9 is X simply add 1 1/3 to 7 and the 2 coordinate points would be 0,1 1/3 and 9,8 1/3.

OpenStudy (triciaal):

This will always work and it's fast. rewrite in the form y = first find the boundary when = the points to use when x = 0 and when y = 0 pick a value of x and see where y lies in relation to the desired inequality

OpenStudy (triciaal):

7x − 9y ≥ 12 y =-7/9 x - 12/9 when x = 0 y = -4/3 when y = 0 x = 12/7 you have 2 points now draw the line when = pick an x value and check the inequality y>

OpenStudy (triciaal):

@math425 do you need more help or can you do the graph now?

OpenStudy (anonymous):

I think I have it. I really appreciate you taking the time to help. Thank you!

hero (hero):

Let me write up something just to make it even clearer

hero (hero):

1. First find the x and y intercepts: ( I know @trichaal provided this) When x = 0, y = -4/3 When y = 0, x = 12/7 2. Plot the points (0, -4/3) and (12/7, 0) https://www.desmos.com/calculator/mhq4hsncnh 3. Draw the line 7x - 9y = 12 https://www.desmos.com/calculator/bmflj1hyzm 4. Test the point (0,0) If true, then shade the part that includes the origin If false, shade the part that doesn't include the origin 7(0) - 9(0) ≥ 12 0 ≥ 12 False 5. Shade the part of the xy plane to the right of 7x - 9y ≥ 12: https://www.desmos.com/calculator/zlyzbsnb8v

hero (hero):

@math425

OpenStudy (anonymous):

okay got it thanks so much.